OpenStudy (anonymous):

Physics: Linear thermal expansion of aircraft carrier in the Mediterranean Sea. A steel aircraft carrier is 370m long when moving through the icy North Atlantic at a temperature of 2.0 degrees Celsius. By how much does the carrier lengthen when it is traveling in the warm Mediterranean Sea at a temperature of 21 degrees Celsius

OpenStudy (anonymous):

and also this one.... A test tube contains 2.54x10-4 m3 ofliquid carbon tetrachloride at a temperature of 75.0°C. The testtube and the carbon tetrachloride are cooled to a temperature of-13.0°C, which is above the freezing point of carbontetrachloride. Find the volume of carbon tetrachloride in the testtube at -13.0°C.

OpenStudy (anonymous):

Well, if you know something called the co-efficient of linear expansion of steel (alpha), you can find out the change in length by using \[DeltaL=L _{0}\alpha DeltaT\]

OpenStudy (anonymous):

All credit goes to Matthias: dl = l0 * a * dT With dl : change of length (delta length, in metres) l0 : original length (in metres) a : Thermal expansion coefficient (for steel it is 11-12 in 10^-6 1/K) dT : difference of temperature, here 19 K (21 - 2 Kelvin), so the carrier lengthen dl = 370m * 11.5E-6 1/K * 19K = 0.080845m ~ 8cm Source: http://answers.yahoo.com/question/index?qid=20101212085656AABVRPC

OpenStudy (anonymous):

for the second problem use this formula: V=Vo(1+B(delta)t)

OpenStudy (anonymous):

wow. your awesome. you've been a great help to me. :)

OpenStudy (anonymous):

No worries, BELIEVE it or not most of the questions people post on here are all over the internet. Just do a quick search if you have any more questions and im sure you'll find the answer to it!